# Provably Efficient Multi-Objective Bandit Algorithms under Preference-Centric Customization

### v1

@@ 6.3 Theoretical Results
Under Assumption 6.2, for any $10<\alpha<$ 1, $\omega>0$, with a probability at least $1-\vartheta$, the preference estimator $\hat{c}_{t}$ in Algorithm 2 verifies for all $t\in[1,T]$:
$\|\hat{\boldsymbol{c}}_{t}-\boldsymbol{\overline{c}}\|_{\boldsymbol{V}_{t-1}}\leq R\sqrt{\omega D\log\big{(}(1+\omega t/\lambda)/\vartheta\big{)}}+\sqrt{\lambda}.$
Please see Appendix D.2 for the proof. This estimate confidence bound essentially implies the effectiveness of WLS preference estimator with the designed weight to handle the random-mapping issue in our problem. With this in hand, we can then derive an upper confidence bound for the expected overall reward.
@@
Set $\beta_{t}=\sqrt{\omega D\log\big{(}(1+\omega t/\lambda)/\vartheta\big{)}}+\sqrt{\lambda}$, then for any $i\in[K]$ and any $t>0$, with probability at least equal to $1-\vartheta-D\alpha^{2}/t^{2}$, we have
$\boldsymbol{\overline{c}}^{\top}\boldsymbol{\mu}_{i}\leq\boldsymbol{\hat{c}}_{t}^{\top}\boldsymbol{\hat{r}}_{i,t}+B_{i,t}^{r}+B_{i,t}^{c},$ (11)
with $B_{i,t}^{r}=\rho_{i,t}^{\alpha}\|\hat{\boldsymbol{c}}_{t}\|_{1}$ and $B_{i,t}^{c}=\beta_{t}\left\|\hat{\boldsymbol{r}}_{i,t}\!+\!\rho_{i,t}^{\alpha}\boldsymbol{e}\right\|_{\boldsymbol{V}_{t-1}^{-1}}$ as the bonus terms for dual-exploration.
Please see Appendix D.3 for the proof. Lemma 5 essentially suggests an upper confidence bound of the expected reward, which is adopted in our dual-exploration policy (8). Notably, the two bonus terms strike a balance between local and global explorations while guaranteeing optimization under the principle of optimism in face of uncertainty.
@@
$\textstyle R(T)=$ $O\Big{(}{\color[rgb]{0,1,1}DR\sqrt{\omega T\log^{2}\Big{(}\big{(}1+\omega T/\lambda\big{)}/\vartheta\Big{)}}}+{\color[rgb]{0,.5,.5}\frac{DR}{\sqrt{\lambda}}\sqrt{\omega DKT\log^{2}\left(\big{(}1+\omega T/\lambda\big{)}/\vartheta)\right)}}+{\color[rgb]{1,.5,0}\sqrt{KT\log\left(T/\vartheta\right)}}+M\Big{)},$
@@
with $M=\left\lfloor\min\big{\{}t^{\prime}\mid t\sigma^{2}_{r\downarrow}+\lambda\geq 2D\omega\sqrt{Kt\log\frac{t}{\alpha}},\forall t\geq t^{\prime}\big{\}}\right\rfloor$ [^2] Since $\sigma^{2}_{\boldsymbol{r}\downarrow}\!\in\!\mathbb{R}^{+}$, $\lim_{t\rightarrow\infty}2D\omega\sqrt{Kt\log\frac{t}{\alpha}}/\big{(}\sigma^{2}_{r\downarrow}t\big{)}=\lim_{t\rightarrow\infty}C_{1}\sqrt{(\log t-C_{2})/t}=0$, as $\sqrt{\log t}$ grows much slowly compared to $\sqrt{t}$. Hence $M$ exists for sufficiently large $t^{\prime}$.. Here the first term represents regret from preference estimation error, the third from reward estimation error, and the second from the combined error of both.
@@
Please see Appendix D.5 for the proof. The key difficulty of the proof is to upper-bound the accumulative preference bonus $\sum_{t}B_{i,t}^{c}$. Specifically, we need to quantify the weighted $\ell_{2}$-norm of the empirical estimation $\boldsymbol{\hat{r}}_{a_{t},t}$ with weighting matrix $\boldsymbol{V}_{t}$ constructed by the true reward $\boldsymbol{r}_{a_{t},t}$ instead. This inconsistency renders the classical induction method [19] for deriving $\log(\frac{\det\boldsymbol{V}_{T}}{\det\boldsymbol{V}_{0}})$ infeasible for upper-bounding $\sum_{t}\|\boldsymbol{\hat{r}}_{a_{t},t}\|_{\boldsymbol{V}_{t-1}^{-1}}^{2}$. To resolve this, we first transfer $\|\boldsymbol{\hat{r}}_{a_{t},t}\|_{\boldsymbol{V}_{t-1}^{-1}}$ to $\|\boldsymbol{\mu}_{a_{t}}\|_{\boldsymbol{V}_{t-1}^{-1}}$. Then we show that for sufficiently large $t$, $a\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}[\boldsymbol{V}_{t-1}]^{-1}}$ serves as an upper bound for $\|\boldsymbol{\mu}_{a_{t}}\|_{\boldsymbol{V}_{t-1}^{-1}}$ with constant $a>1$ (see Lemma 15). This allows us to use $a\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}[\boldsymbol{V}_{t-1}]^{-1}}$ as an upper-bound, where a new recursion relationship between $\mathbb{E}[\boldsymbol{V}_{t-1}]$ and $\boldsymbol{\mu}_{a_{t}}$ can be guaranteed, enabling us to bound $\log(\frac{\det\mathbb{E}[\boldsymbol{V}_{T}]}{\det\mathbb{E}[\boldsymbol{V}_{0}]})$ via induction, which can further be bounded by slightly modifying existing techniques in linear bandits.
Remark 6.1.
@@
Theorem 6 shows that, even without explicit preference feedback, PRUCB-HP achieves sub-linear regret through carefully designed mechanisms for preference adaptation. In particular, for $t\geq M$, where $M$ is a constant independent of $T$, the regret asymptotically scales as $\tilde{O}(D\sqrt{T})$. To the best of our knowledge, this is the first result characterizing the performance of PAMO-MAB with hidden preference in the literature.
(a) Unknown preference case
(b) Hidden preference case
Figure 5: Regret comparison of our proposed PRUCB with other benchmarks under different preference environments, where our methods outperforms other methods significantly.
@@ 7 Numerical Analysis
We evaluate the performance of PRUCB-UP and PRUCB-HP in unknown and hidden preference environments, respectively. The PAMO-MAB instance includes $K$ arms and $D$ objectives, with preferences and rewards following Gaussian distributions with randomly initialized means.. (Detailed settings refer to Appendix A.1). For the hidden preference case, we introduce a user-switching protocol to simulate practical scenarios. The environment features multiple users, each exposed to a block of arms (5 in our setup) per round. Only arms within the current block can be selected for that user. In the next round, the arm block rotates to a another user. The objective is to maximize cumulative overall ratings across all users, and performance is measured by the averaged users’ regrets. A more detailed illustration is provided in Appendix A.2.
We compare our results with other baselines including S-UCB, Pareto-UCB [1], S-MOSS, Pareto-TS [13]), UCB [17], MOSS [23] and OFUL [19]. The regret is averaged across 10 trials with round $T=5000$. Figure 5 shows that our proposed algorithms significantly outperform other competitors under both environments. It is worth noting that for all the preference-free competitors exhibit linear regret, aligning with Proposition 1, demonstrating that approaches agnostic to user preferences cannot align their outputs with user preferences, even if they achieve Pareto optimality. For more comprehensive experimental analyses, please refer to Appendix A.
@@ 8 Conclusion
Appendix A Experiments
@@
A.1 Experiments in Unknown Preference Environment
In this section, we verify the capability of PRUCB and PRUCB-UP to model user preference $\boldsymbol{c}_{t}$ and optimize the overall reward in a stationary preference environment. We compare these two algorithms in terms of regret with the following multi-objective bandits algorithms.
  • S-UCB [1]: the scalarized UCB algorithm, which scalarizes the multi-dimensional reward by assigning weights to each objective and then employs the single objective UCB algorithm [17]. Throughout the experiments, we assign each objective with equal weight.
  • S-MOSS: the scalarized UCB algorithm, which follows the similar way with S-UCB by scalarizing the multi-dimensional reward into a single one, but uses MOSS [23] policy for arm selection.
  • Pareto-UCB [1]: the Pareto-based algorithm, which compares different arms by the upper confidence bounds of their expected multi-dimensional reward by Pareto order and pulls an arm uniformly from the approximate Pareto front.
  • Pareto-TS [13]: the Pareto-based algorithm, which makes use of the Thompson sampling technique to estimate the expected reward for every arm and selects an arm uniformly at random from the estimated Pareto front.
  • GT: The PRUCB-UP variant with known preference as a ground truth, where we replace the preference estimation in Algorithm 1 with the true preference estimation.
Experimental settings. For evaluation, we use a synthetic dataset. Specifically, we consider the MO-MAB with $K$ arms, each arm $i\in[K]$ associated with a $D$-dimensional reward, where the reward of each objective $d$ follows a Gaussian distribution with a randomized mean $\boldsymbol{\mu}_{i}(d)\in[0,1]$ and variance of 0.01. For user preference, we consider two settings including predefined preference and randomized preference. For predefined preference-aware structure, we define the mean preference $\boldsymbol{\overline{c}}$ as $\boldsymbol{\overline{c}}(d)=2.0\text{ if }d=j;0.5\text{ otherwise}$, where $j\in[D]$ is randomly selected. The practical implication of this structure is that it represents a common scenario in which the user exhibits a markedly higher preference for one particular objective while showing little interest in others. For randomized preference, the values of mean preference $\boldsymbol{\overline{c}}$ are randomly defined within $[0,5]$. For both setups, the instantaneous preference is generated under Gaussian distributions with corresponding means and variance of 0.5.
Implementations. For the implementations of the algorithms, following previous studies [17, 24], we set $\alpha=1$. The time horizon is set to $T=5000$ rounds, and we repeat 10 trials for each set of evaluation due to the randomness from both environment and algorithms.
Results. We report the averaged regret performance of the algorithms in Fig. 6. It is evident that our algorithm significantly outperform other competitors in all experiments. This is expected since the competing algorithms are designed for Pareto-optimality identification and do not utilize the preference structure of users considered in this paper, which our algorithm explicitly exploits. Additionally, from the zoom-in window, we observe that PRUCB-UP exhibits only a very slight performance degradation compared to GT, which knows the preference expectation in advance. This indicates that the proposed PRUCB-UP can effectively model user preference via empirical estimation in stationary preference environments.
A.2 Experiments in Hidden Preferences Environment

### v2

@@ 4.3 Theoretical Results
Under Assumption 5, for any user $n\in[N]$, $0<\alpha<1$, $\omega>0$, with a probability at least $1-\vartheta$, the preference estimator $\hat{c}_{t}^{n}$ in Algorithm 1 verifies for all $t\in[1,T]$:
$\|\hat{\boldsymbol{c}}_{t}^{n}-\boldsymbol{\overline{c}}^{n}\|_{\boldsymbol{V}_{t-1}^{n}}\leq R\sqrt{\omega D\log\big((1+\omega t/\lambda)/\vartheta\big)}+\sqrt{\lambda}.$
Please see Appendix D.2 for the proof. This estimate confidence bound essentially implies the effectiveness of WLS preference estimator with the designed weight to handle the random-mapping issue in our problem. With this in hand, we can then derive an upper confidence bound for the expected overall reward for each user.
@@
Set $\beta_{t}=\sqrt{\omega D\log\big((1+\omega t/\lambda)/\vartheta\big)}+\sqrt{\lambda}$, then for any $n\in[N]$, $i\in[K]$ and $t>0$, with probability at least equal to $1-\vartheta-D\alpha^{2}/t^{2}$, we have
${\boldsymbol{\overline{c}}^{n}}^{\top}\boldsymbol{\mu}_{i}\leq{\boldsymbol{\hat{c}}_{t}^{n}}^{\top}\boldsymbol{\hat{r}}_{i,t}+B_{i,t}^{n,r}+B_{i,t}^{n,c},$ (8)
with $B_{i,t}^{n,r}\!=\!\rho_{i,t}^{\alpha}\|\hat{\boldsymbol{c}}_{t}^{n}\|_{1}$ and $B_{i,t}^{n,c}\!=\!\beta_{t}\left\|\hat{\boldsymbol{r}}_{i,t}\!+\!\rho_{i,t}^{\alpha}\boldsymbol{e}\right\|_{{(\boldsymbol{V}_{t-1}^{n})}^{-1}}$ as the bonus terms for dual-exploration.
Please see Appendix D.3 for the proof. Lemma 3 essentially suggests an upper confidence bound of the expected reward for each user, as adopted in our dual-exploration policy (Eq. 5). Notably, two bonus terms strike a balance between local and global explorations while guaranteeing optimization under the principle of optimism in face of uncertainty.
@@
$\textstyle R(T)=$ $O\Big(DRN\sqrt{\omega T\log^{2}\Big(\big(1+\omega T/\lambda\big)/\vartheta\Big)}$
@@
$\qquad+\frac{DRN}{\sqrt{\lambda}}\sqrt{\omega DKT\log^{2}\left(\big(1+\omega T/\lambda\big)/\vartheta)\right)}$
$\qquad+N\sqrt{KT\log\left(T/\vartheta\right)}+MN\Big),$
@@
with $M=\left\lfloor\min\big\{t^{\prime}\mid t\sigma^{2}_{r\downarrow}+\lambda\geq 2D\omega\sqrt{Kt\log\frac{t}{\alpha}},\forall t\geq t^{\prime}\big\}\right\rfloor$. [^2] Since $\sigma^{2}_{\boldsymbol{r}\downarrow}\!\in\!\mathbb{R}^{+}$, $\lim_{t\rightarrow\infty}2D\omega\sqrt{Kt\log\frac{t}{\alpha}}/\big(\sigma^{2}_{r\downarrow}t\big)=\lim_{t\rightarrow\infty}C_{1}\sqrt{(\log t-C_{2})/t}=0$, as $\sqrt{\log t}$ grows much slowly compared to $\sqrt{t}$. Hence $M$ exists for sufficiently large $t^{\prime}$. Here the first term represents regret from preference estimation error, the third from reward estimation error, and the second from the combined error of both.
@@
Please see Appendix D.5 for the proof. The key difficulty of the proof is to upper-bound the accumulative preference bonus $\sum_{t}B_{i,t}^{c}$. Specifically, we need to quantify the weighted $\ell_{2}$-norm of the empirical estimation $\boldsymbol{\hat{r}}_{a_{t},t}$ with weighting matrix $\boldsymbol{V}_{t}$ constructed by the true reward $\boldsymbol{r}_{a_{t},t}$ instead. This inconsistency renders the classical induction method (Abbasi-Yadkori, Pál, and Szepesvári 2011) for deriving $\log(\frac{\det\boldsymbol{V}_{T}}{\det\boldsymbol{V}_{0}})$ infeasible for upper-bounding $\sum_{t}\|\boldsymbol{\hat{r}}_{a_{t},t}\|_{\boldsymbol{V}_{t-1}^{-1}}^{2}$. To resolve this, we first transfer $\|\boldsymbol{\hat{r}}_{a_{t},t}\|_{\boldsymbol{V}_{t-1}^{-1}}$ to $\|\boldsymbol{\mu}_{a_{t}}\|_{\boldsymbol{V}_{t-1}^{-1}}$. Then we show that for sufficiently large $t$, $a\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}[\boldsymbol{V}_{t-1}]^{-1}}$ serves as an upper bound for $\|\boldsymbol{\mu}_{a_{t}}\|_{\boldsymbol{V}_{t-1}^{-1}}$ with constant $a>1$ (see Lemma 8). This allows us to use $a\|\boldsymbol{\mu}_{a_{t}}\|_{\mathbb{E}[\boldsymbol{V}_{t-1}]^{-1}}$ as an upper-bound, where a new recursion relationship between $\mathbb{E}[\boldsymbol{V}_{t-1}]$ and $\boldsymbol{\mu}_{a_{t}}$ can be guaranteed, enabling us to bound $\log(\frac{\det\mathbb{E}[\boldsymbol{V}_{T}]}{\det\mathbb{E}[\boldsymbol{V}_{0}]})$ via induction, which can further be bounded by slightly modifying existing techniques in linear bandits.
Remark 1.
@@
Theorem 1 shows that, even without explicit preference feedback, PRUCB-HP achieves sub-linear regret through carefully designed mechanisms for preference adaptation. In particular, for $t\geq M$, where $M$ is a constant independent of $T$, the regret asymptotically scales as $\tilde{O}(D\sqrt{T})$. To the best of our knowledge, this is the first result characterizing the performance of PAMO-MAB with hidden preference.
5 A Special Case with Provided Preference
We next consider another practically important case where the user’s preferences are explicitly provided to the agent either before or after decision making. This setup is prevalent in numerous real-world applications. Many online systems now allow users to express their preferences through interactive techniques, such as conversations and prompt design, either before or after taking action. For example, as shown in Fig. 4, a user can explicitly share her movie preferences with an LLM-based chat system either prior to receiving a recommendation ( in the initial input) or after receiving one ( in the follow-up conversation).
We observe that when the preference vector $\boldsymbol{c}t$ is known before decision-making, the problem reduces to a standard single-objective MAB. Instead, we focus on the more challenging case where the preference is unknown during action selection but revealed afterward. The known-before-action case can be treated as a special instance of this setting and is deferred to Appendix E.2. The goal remains to maximize the cumulative overall-reward over all users. Formally, at each round $t$, the learner selects an arm $a_{t}$, then observes both the reward $\boldsymbol{r}\_{a_{t},t}$ and user’s preference $\boldsymbol{c}_{t}$, with both adhering to the original Assumptions 1 and 2. This setting avoids the challenges of preference inference and stochastic reward mapping, and can be viewed as a special case of the hidden preference setting with zero noise in preference feedback.
To address this case, we propose a simplified algorithm: PRUCB-UP (Algorithm 2) which builds on the same Preference-Aware framework used in the hidden case, with two direct adaptations:
Preference estimation.
Due to the explicitly provided preference feedback $\boldsymbol{c}_{t}^{n}$, we can directly leverage the empirical average of historical feedback as the preference estimate for each user $n$. For $t\geq 1$, preference estimate is updated as
$\hat{\boldsymbol{c}}^{n}_{t+1}=\big((t-1)\boldsymbol{\hat{c}}^{n}_{t}+\boldsymbol{c}_{t}^{n}\big)/t.$ (9)
$N_{i,t+1}=N_{i,t}+\sum_{n\in[N]}\mathds{1}_{\{a_{t}^{n}=i\}},$ (10)
$\hat{\boldsymbol{r}}_{i,t+1}=(\hat{\boldsymbol{r}}_{i,t}N_{i,t}+\sum_{n\in[N]}\boldsymbol{r}_{a_{t}^{n},t}\cdot\mathds{1}_{\{a_{t}^{n}=i\}})/N_{i,t+1},$
Preference-aware optimization.
Algorithm 2 PRUCB Unknown Preference (PRUCB-UP)
Initialize: $\alpha$, $N_{i,1}\!\leftarrow\!0$, $\boldsymbol{\hat{r}}_{i,1}\!\leftarrow\![0]^{D},\forall i\!\in\![K]$; $\quad\quad\quad\quad\boldsymbol{\hat{c}}_{1}^{n}\!\leftarrow\![0]^{D},\forall n\in[N]$.
for $t=1,\cdots,T$ do
  for user $n\in[N]$ do
   Draw arm $a_{t}^{n}$ by (11),
   Observe reward $\boldsymbol{r}_{a_{t}^{n},t}$ and user preference $\boldsymbol{c}_{t}^{n}$.
   Update preference estimate $\hat{\boldsymbol{c}}^{n}_{t+1}$ by (9).
  end for
  Update $N_{i,t+1}$, $\hat{\boldsymbol{r}}_{i,t+1},\forall i\in[K]$ by (10).
end for
We adopt the same UCB framework as in hidden preference case for preference-aware optimization, but with key simplifications. Since the preference $\boldsymbol{c}_{t}$ is observed after action selection and is independent of the chosen arm, its estimation does not involve sequential decision-making—hence no confidence term is needed for $\hat{\boldsymbol{c}}_{t}$. In contrast, reward estimates $\hat{\boldsymbol{r}}_{t}$ remain action-dependent, requiring reward bonuses to ensure sufficient exploration.
To this end, the arm selection policy is designed as:
$a_{t}^{n}=\operatorname*{arg\,max}_{i\in\mathcal{A}^{n}_{t}}({\hat{\boldsymbol{c}}_{t}^{n}})^{\top}(\hat{\boldsymbol{r}}_{i,t}+\rho_{i,t}^{\alpha}\boldsymbol{e}).$ (11)
where $\rho_{i,t}^{\alpha}=\sqrt{\log(t/\alpha)/\max\{1,N_{i,t}\}}$ is standard Hoeffding bonus. We characterize the regret of PRUCB-UP below.
Theorem 2.
Let $\mathcal{T}_{i}^{n}\!=\!\{t\!\in\![T]\mid i\!\in\!\mathcal{A}_{t}^{n},i\!\neq\!a_{t}^{n*}\}$, $\eta_{i}^{n\uparrow}=\max_{t\in\mathcal{T}_{i}^{n}}{\overline{\boldsymbol{c}}^{n}}^{\top}(\boldsymbol{\mu}_{a_{t}^{n*}}-\boldsymbol{\mu}_{i})$, $\eta_{i}^{n\downarrow}=\min_{t\in\mathcal{T}_{i}^{n}}{\overline{\boldsymbol{c}}^{n}}^{\top}(\boldsymbol{\mu}_{a_{t}^{n*}}-\boldsymbol{\mu}_{i})$, $\|\Delta_{i}^{n\uparrow}\|_{2}=\max_{\{t,j\}\in\mathcal{T}_{i}^{n}\times\mathcal{A}^{n}_{t}}\|\boldsymbol{\mu}_{i}-\boldsymbol{\mu}_{j}\|_{2}$, Algorithm 2 has the regret R(T) of
$O\Bigg(\sum_{n\in[N]}\sum_{i\in[K]}\Big(\underbrace{\frac{\delta^{2}\log T}{\eta_{i}^{n\downarrow}}+D\pi^{2}\alpha^{2}\eta_{i}^{n\uparrow}}_{R^{r}_{T}}+\underbrace{\frac{D^{2}\|\Delta_{i}^{n\uparrow}\|_{2}^{2}\delta^{2}}{\eta_{i}^{n\downarrow}}}_{R^{c}_{T}}\Big)\Bigg),$
Remark 2.
Theorem 2 shows that with preference feedback, PRUCB-UP achieves a regret of $O(KN\delta\log T+KN\delta D^{2})$, demonstrating near-optimal performance. Notably, the regret caused by additional preference estimation error is bounded by a constant related to objective dimension $D$ and preference $\ell_{1}$-norm bound $\delta$. This implies that the impact of preference estimation error on the regret is small. Additional, we show the preference known case can be solved by PRUCB-UP as a special case, achieving regret of $O(K\delta\log T)$, please see Appendix E.2 for details.
To prove Theorem 2, the main difficulty lies in decoupling and capturing the effect of the joint error from both reward estimation and preference estimation on the final regret. To address this, we introduce a tunable parameter $\epsilon$ to quantify the error of preference estimate $\boldsymbol{\hat{c}}_{t}$, and decompose suboptimal actions into two disjoint sets: (1) suboptimal pulls under sufficiently precise preference estimate; (2) suboptimal pulls under imprecise preference estimate, which we show to be more analytically tractable. Please see Appendix E.1 for the full proof.
@@ 6 Numerical Analysis
We evaluate the performance of PRUCB-UP and PRUCB-HP in unknown and hidden preference environments, respectively. The PAMO-MAB instance includes $N$ users, $K$ arms and $D$ objectives, with preferences and rewards following Gaussian distributions with randomly initialized means. (Detailed settings refer to Appendix B.2). We introduce a user-switching protocol to simulate practical scenarios. The environment features multiple users, each exposed to a block of arms (5 in our setup) per round. Only arms within the current block can be selected for that user. In the next round, the arm block rotates to a another user. The objective is to maximize cumulative overall ratings across all users. A more detailed illustration is provided in Appendix B.2.
(a) Regret comparison of our proposed PRUCB with other benchmarks under different preference environments, where our methods outperforms other methods significantly.
We compare our results with other baselines including S-UCB, Pareto-UCB (Drugan and Nowe 2013), S-MOSS, Pareto-TS (Yahyaa and Manderick 2015)), UCB (Auer, Cesa-Bianchi, and Fischer 2002), MOSS (Audibert and Bubeck 2009) and OFUL (Abbasi-Yadkori, Pál, and Szepesvári 2011). The regret is averaged across 10 trials with round $T=5000$. Figure 6(a) shows that our proposed algorithms significantly outperform other competitors under both environments. It is worth noting that for all the preference-free competitors exhibit linear regret, aligning with Proposition 1, demonstrating that approaches agnostic to user preferences cannot align their outputs with user preferences, even if they achieve Pareto optimality. For more comprehensive experimental analyses, please refer to Appendix B.
@@ 7 Conclusion
Appendix A Related Work
Multi-Objective Multi-Armed Bandits.
MO-MAB extends scalar rewards in the standard MAB problem to multi-dimensional vectors. The Pareto-UCB (Drugan and Nowe 2013) introduced the MO-MAB framework and Pareto regret as a metric, achieving $O(\log T)$ Pareto regret using the UCB technique. Other techniques, including Knowledge Gradient (Yahyaa, Drugan, and Manderick 2014) and Thompson Sampling (Yahyaa and Manderick 2015), have subsequently been adapted for MO-MAB. Additionally, researchers have extended the contextual setup to MO-MAB (Turgay, Oner, and Tekin 2018; Lu et al. 2019). These studies aim to approximate the entire Pareto front $\mathcal{O}^{*}$, and employ a random arm selection policy on the estimated Pareto front for Pareto optimality. Considering computing the full Pareto front is expensive, another line of work propose to converts the multi-dimensional reward into a scalar value through a scalarization function, targeting a specific Pareto optimal solution while not the entire Pareto front. Our work also falls within the realm of this framework. The scalarization function can either be randomly initialized (chosen) (Drugan and Nowe 2013; Xu and Klabjan 2023), or optimized based on a fixed metric, such as the Generalized Gini Index score (Busa-Fekete et al. 2017; Mehrotra, Xue, and Lalmas 2020). Nonetheless, existing studies primarily achieve Pareto optimality through a global policy for arm selection across all users. As discussed in Section 1, merely achieving Pareto optimality with a global policy may not yield favorable outcomes, as users have diverse preferences on different objectives.
Preference-based MO-MAB optimization.
Recent studies have explored MO-MAB optimization using lexicographic order (Ehrgott 2005) to reflect user preferences. In lexicographic order, objectives are prioritized hierarchically, where the first objective takes absolute precedence over the second, and so on. Hüyük and Tekin (2021) first introduced lexicographic order to MO-MAB, and Cheng et al. (2024) extended it to mixed Pareto-lexicographic environments. However, lexicographic order may not adequately capture a user’s overall satisfaction in real-world applications, where preferences often involve trade-offs rather than strict prioritization. For example, a user may prefer a \$10 meal with good taste over a \$9.5 meal with poor taste, even though cost is a priority. Our work proposes a more general framework that incorporates a weighted order based on the user’s explicit preference space. Notably, the lexicographic order becomes a special case of our proposed PAMO-MAB framework.
Appendix B Experiments
@@
B.1 Experiments in Hidden Preferences Environment